2g of NaOH is dissolved in 500dm³ of solvent.what is the molar concentration of the solution?

1
Expert's answer
2022-03-17T15:01:39-0400

Solution:

The molar mass of NaOH is 40 g/mol

Therefore,

Moles of NaOH = (2 g NaOH) × (1 mol NaOH / 40 g NaOH) = 0.05 mol NaOH


Convert dm3 to L:

1 dm3 = 1 L

(500 dm3) × (1 L / 1 dm3) = 500 L

Liters of solution = 500 L


Molar concentration = Moles of solute / Liters of solution

Therefore,

Molar concentration of NaOH solution = (0.05 mol) / (500 L) = 0.0001 mol/L= 1×10−4 M

Molar concentration of NaOH solution = 1×10−4 M


Answer: The molar concentration of the solution is 1×10−4 M